OpenStudy (anonymous):

Solve: -1/2x + 10 = −x + 12

OpenStudy (maheshmeghwal9):

take variable terms one side & constant terms one side & then do simple arithmetic of variables & constants on respective sides. then u will get 'x' which is variable.

OpenStudy (anonymous):

8

OpenStudy (anonymous):

-1/2x + 10 = −x + 12 now take LCM -1+20x ------- = -x+12 2x -1+20x=2x(-x+12) -1+20x=-2x^2+24x 2x^2-24x+20x-1=0 2x^2-4x-1=0 now for further calculation use quadratic equation where a=2, b=-4 and c=-1 Hope it helps :)

OpenStudy (anonymous):

is it -4

OpenStudy (anonymous):

what the answer?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

no i am asking that are you saying the answer is -4 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the answer -4